# Quantifying the Robustness of Retrieval-Augmented Language Models Against Spurious Features in Grounding Data

### v1

@@ 3.4 Causal Features Preservation
To eliminate the effect of causal features, it is essential to follow the principle of controlled experiments by keeping causal features constant while systematically manipulating spurious features. This approach isolates the impact of spurious features from that of causal features, enabling an accurate quantification of robustness against spurious features. In our framework, we introduce two methods to ensure the stability of causal features in the grounding data.
@@
In general, this algorithm can be implemented by any NLI system. However, in our case, the concatenation of $G$ and $g(G)$ sometimes exceeds the context limitation of a Bert-based NLI model. Hence, we apply an LLM-based NLI system  [^4] Farquhar et al. (2024) confirms the effectiveness of the LLM-based NLI system through human annotation, demonstrating that its performance is on par with the DeBERTa-large model used in Kuhn et al. (2023). to implement the bidirectional entailment algorithm. The prompt is included in Appendix D.
@@
While semantic equivalence protects causal features to the greatest extent, the perturbation may lead to the correct answer being paraphrased into an alias (e.g., ”President Roosevelt” to ”Roosevelt”). These variations in the grounding data are likely to result in false negatives when determining response correctness, despite the NQ-Open dataset providing multiple potential answer variants for each query. To address this issue, we employ a simple string-matching strategy to filter out documents that undergo unexpected modifications. For Golden documents that originally contained the correct answers, we keep them only if they preserve the ground truths after perturbation. For Noise documents that did not contain the correct answers, we discard them if they unexpectedly acquire the ground truths due to perturbations.
3.5 Overview of the Synthetic Dataset
Through the steps of spurious features injection and causal features preservation, we derive the final dataset available for robustness evaluation. The synthetic dataset generated by the SURE framework is divided into four subsets based on the categories of queries and documents within the instances. Notably, the distribution of the dataset is model-specific, as the classification of Known and Unknown queries is determined by the intrinsic knowledge of the target LLM. Table 1 presents the dataset statistics for assessing Mistral-7B-Instruct-v0.3. And the distribution for Llama-3.1-8B-Instruct is shown in Appendix E.
@@ 3.5 Overview of the Synthetic Dataset
Table 1: Statistics of the evaluation dataset for Mistral-7B-Instruct-v0.3. K-G denotes the instances composed of (Known query, Golden Document), while U-N refers to the instances consisting of (Unknown query, Noise Document). The values represents the number of instance pairs for each type of perturbations within the category of spurious features.
3.6 Robustness Evaluation
To quantify whether a RAG system is robust and unbiased at the instance level, we assign a ternary label to each instance by comparing the correctness of the LLM’s response before and after introducing the perturbation. This comparison process can be formulated as $C=Y(y_{i})-Y(\hat{y_{i}})$, where $C$ lies in the set $(-1,0,1)$. Based on the comparison outcomes, we define three metrics: Robustness Rate (RR), Win Rate (WR), and Lose Rate (LR). The RR is calculated as follows:
$\text{RR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(C==0)$ (5)
where $N$ is the total number of instances in the dataset; $y_{i}$ and $\hat{y_{i}}$ represent the outputs of LLM for the original and perturbed instances. RR measures the proportion of instances where the RALM’s answer remains consistent (0) before and after introducing the perturbation. Similarly, WR and LR quantify the proportions of instances where the correctness of the RALM’s response changes after the perturbation, either from incorrect to correct ($C==-1$) or from correct to incorrect ($C==1$).
Figure 3: A comparison of dataset-level metric (Acc) and instance-level metric (RR) for robustness evaluation. ✔and ✗indicate the correctness of responses. In this example, the value of RR reflects the unrobustness at the instance level, while dataset-level metric overlook the sensitivity of RALMs to spurious features within documents.
4 Experiments
Mistral-7B-Instruct-v0.3
Spurious Features Perturbations Known-Golden Known-Noise Unknown-Golden Unknown-Noise Total
LR RR WR LR RR WR LR RR WR LR RR WR LR RR WR
Style Simple 7.33 85.00 7,67 4.45 91.64 3.90 7.87 82.95 9.18 0.70 98.76 0.54 3.06 94.09 2.85
Complex 6.05 87.42 6.53 3.85 92.03 4.12 6.90 85.92 7.17 0.62 98.82 0.56 2.63 94.61 2.76
@@ 4 Experiments
Self-Generated 6.30 87.06 6.64 3.94 92.02 4.04 6.26 86.80 6.94 0.61 98.77 0.62 2.74 94.42 2.85
[…]
Random 4.47 91.87 3.66 2.43 95.15 2.42 4.18 91.44 4.38 0.36 99.27 0.37 1.76 96.56 1.68
LLM-Ranked 3.52 93.15 3.33 2.07 95.84 2.09 3.57 92.89 3.54 0.34 99.30 0.36 1.48 97.04 1.48
[…]
HTML 9.30 87.03 3.67 5.89 92.36 1.74 8.36 87.39 4.25 0.74 99.01 0.26 4.00 94.62 1.38
YAML 4.75 90.90 4.35 3.88 93.24 2.87 5.05 90.53 4.42 0.51 99.06 0.44 2.47 95.55 1.98
Markdown 3.98 92.49 3.53 2.91 94.36 2.72 4.11 92.59 3.31 0.44 99.15 0.41 1.94 96.29 1.77
[…]
Timestamp (post) 2.74 94.87 2.40 1.16 97.63 1.21 3.45 94.41 2.14 0.17 99.68 0.15 0.98 98.12 0.89
Datasource (wiki) 3.78 92.31 3.91 1.5 96.66 1.84 3.69 92.95 3.36 0.26 99.48 0.26 1.28 97.31 1.41
Datasource (twitter) 2.68 93.59 3.73 1.3 97.22 1.48 2.04 94.90 3.06 0.20 99.59 0.21 0.98 97.80 1.22
Table 2: Robustness evaluation results of Mistral-7B-Instruct-v0.3 on the synthetic dataset. For timestamp perturbation, (pre) and (post) represents whether the virtual timestamp we created is before or after the LLM’s knowledge cutoff date. For datasource perturbation, (wiki) and (twitter) indicate the domains of the pseudo link we synthesize. We use Bold to mark the WR values that are higher than the LR, suggesting that the perturbation is beneficial.
4.1 Experimental Setup
We assess the robustness of RAG systems to spurious features by evaluating them on their most popular application—the Question Answering (QA) task, following the standard ”retrieve-read” setting of the RAG paradigm. We employ Contriever-msmarco as the default retriever and conduct our main experiments using the synthetic dataset generated by the SURE framework. Due to space limitations, additional implementation details can be found in Appendix F.
4.2 Result Analysis
In this section, we analyze our main results by category and offer insights from various perspectives. The results of Mistral-7B-Instruct and Llama-3.1-8B-Instruct are presented in Table 2 and Table 6, respectively.
For Different Queries
There is a significant difference in robustness rates between known and unknown queries when evaluated on noise documents (compare K-N and U-N), whereas no such gap is observed for golden documents (compare K-G and U-G). This phenomenon arises from the tug-of-war between an LLM’s internal prior and external evidence. When the retrieved content is noise documents, LLMs are likely to override it with their own correct prior knowledge (Known). However, the robustness to explicit noise may be affected by the spurious features in documents. In other words, implicit noise (spurious features) can coexist with explicit noise, highlighting the challenge of maintaining robustness to spurious features in complex noise scenarios.
For Different Grounding Data
Across all types of spurious features, the robustness rate on noise documents is consistently higher than that on golden documents, whether for known or unknown queries. This derives from the limitation of the RR metric we propose, which measures unrobustness by capturing changes in answer correctness rather than the inconsistency of responses. When tested on noise documents, regardless of which spurious features are injected, LLMs always generate incorrect responses, as noise documents lack ground truths. In this case, even though the responses change, the RR does not decrease since all responses remain incorrect. However, in real-world applications, our primary concern is whether the RAG system can consistently generate correct answer when faced with golden documents containing different spurious features. Therefore, we primarily focus on the RR results for the (K, G) and (U, G) subsets.
For Different Metrics
If the win rate surpasses the lose rate, it shows that more instances were corrected rather than misanswered after introducing perturbations. Based on the comparison of WR and LR in Table 2 and Table 6, we can conclude that Not every spurious feature is harmful and they can even be beneficial sometimes. For example, the WR of source perturbations is consistently higher than the LR on both models. This suggests that the performance of RAG can be enhanced by introducing beneficial spurious features (e.g., simply paraphrasing documents using LLMs).
For Different Perturbations
We observe notable differences in the robustness rates among five types of spurious features. However, within each category, the robustness rates across different perturbations are relatively close. When further comparing perturbations within the same category, we find that while their RR values are comparable, their WR and LR can differ significantly, indicating LLMs’ preference for certain sub-perturbations. In summary, there are variations in robustness across different spurious features, and preference exists among the sub-perturbations of each spurious feature.
4.3 SIG Benchmark & Further Analysis
The raw synthetic dataset is not ideal for extensive evaluation due to its large size. Furthermore, the class imbalance result in unfair comparisons across different types of spurious features. To facilitate more efficient evaluation, we extract the most challenging data from the subsets of synthetic dataset to create a lightweight benchmark: SIG (Spurious features In Golden document). Specifically, we select instance pairs where both the Mistral and Llama models exhibit sensitivity and unrobustness in our main experiments. An equal number of samples (100) are chosen from the (K,G) and (U,G) subsets of each perturbations.
Robustness Comparison of SOTA LLMs
We evaluate a broader range of models on SIG benchmark, including GPT-4O, GPT-4O-mini, Mistral-Large-Instruct  [^5] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama-3.3-70B-Instruct, Deepseek-v3 (671B,MoE), and Qwen2.5-72B-Instruct. To better compare the robustness of different models, we average the RR of each perturbation within a category to derive the overall robustness for a specific type of spurious feature. The performance of six SOTA LLMs is then visualized using a radar chart, as shown in Figure 4. Notably, GPT-4O-mini achieved the best performance, even surpassing GPT-4O by a large margin on the format and meta types. Despite the impressive robustness of closed-source models, they may still exhibit sensitivity to certain specific perturbations. For instance, GPT-4o achieved only an 89% robustness rate on the datasource(twitter) perturbation.
Figure 4: Comparison of Robustness rates across six SOTA LLMs.
Scaling Analysis for Different Model Sizes
To investigate the impact of parameter scale on RAG robustness, we gradually increase the size of LLM-based readers (Qwen2.5 series, ranging from 0.5B to 72B) and evaluate their robustness across five types of spurious features. As illustrated in Figure 5, the robustness rate for all spurious features shows a relatively upward trend as the model size increases. However, when we further scale the model from 32B to 72B, the RR undergoes a significant decline (except for format and meta). This indicates that robustness issues related to spurious features cannot be resolved simply by increasing model size. Interestingly, for meta perturbations, while RALMs demonstrate strong robustness across all scales (even for the 0.5B model), their performance receives little to no benefit from scaling up.
Figure 5: Scaling analysis of robustness to spurious features.
5 Conclusion
In this work, we formally highlight the spurious features problem in RAG system. Through preliminary experiments, we provide statistical evidence to support the presence of spurious features in RALMs. We also propose a novel evaluation framework, SURE, to assess the robustness of LLMs against spurious features. This framework includes a comprehensive taxonomy of spurious features, carefully designed metrics for robustness evaluation, and a data synthesis pipeline. Extensive evaluation and further analysis offers valuable insights for future research.
Impact Statement
We introduce a new type of noise, spurious features, that can undermine the robustness and trustworthiness of contemporary RAG systems. With the increasing diversity of golden documents on the internet, this noise may become more prevalent. Furthermore, traditional methods for improving the robustness of RAG systems are ineffective for handling spurious features, as they are implicit and can coexist with golden documents. Therefore, harmful spurious features could undermines user trust and contaminates Internet data. Moving forward, we aim to explore methods for mitigating hallucinations and robustness issues caused by spurious features within grounding data.
Appendix A Related Work
A.1 Robustness Evaluation of Retrieval-Augmented Generation
RAG systems comprise two core components: a retriever and an LLM-based reader. Augmenting LLMs with retrieved external knowledge has been proven to effectively reduce hallucinations (Shuster et al., 2021; Kang et al., 2023). However, the retrieved contexts inevitably contains noise in addition to desirable knowledge, which may mislead LLMs to produce an incorrect response (Bian et al., 2024; Feldman et al., 2024). Previous works have explored automated evaluation frameworks to assess the robustness of RAG systems in various settings. For instance, Chen et al. (2024a) benchmarked four fundamental capabilities required for RAG, including noise robustness, negative rejection, information integration and counterfactual robustness. Some studies have provided a detailed taxonomy of noise documents to further simulate the complexity of real-world scenarios and highlighted the potential positive effects of certain types of noise (Cuconasu et al., 2024; Wu et al., 2024a). There are also some recent works that propose using LLM-as-a-judge (Li et al., 2024a) to evaluate the RAG system (Wang et al., 2024b).
While these studies have identified several explicit noises that affect the robustness of RAG systems, they overlook implicit noises. This type of noise, such as phrasing and formatting, is everywhere and unavoidable, as it coexists with the grounding data without altering its semantic information. In this work, we define these semantic-agnostic noises as spurious features and evaluate the robustness of RALMs to such noises.
A.2 Prompt Sensitivity of LLMs
Large Language Models take prompts as inputs and then generate response accordingly. Prompts are instructions provided to an LLM to perform specific tasks automatically and ensure desired qualities in the generated output. However, it is known that current LLMs are sensitive to the features of input prompts (Zhu et al., 2023). This sensitivity poses challenges for researchers attempting to evaluate the model’s performance accurately and precisely (Zhuo et al., 2024).
Some existing works have investigated the impact of different prompt techniques on model performance, including chain-of-thought (Wei et al., 2022), in-context learning (Min et al., 2022), and role-play prompting (Kong et al., 2024). Beyond these causal features that significantly influence the meaning of prompts, other works have demonstrated that LLMs are highly sensitive to spurious features (Sclar et al., 2024), e.g, prompt formatting (He et al., 2024), language style (Li et al., 2023a), the order of options (Pezeshkpour & Hruschka, 2024).
Currently, there is no statistical or empirical evidence to support the existence of spurious features in RALMs. To address this gap, we extend the definition of spurious features to RAG systems through statistical testing and empirical analysis.
Appendix B Preliminary Experiment Results
We introduce several semantic-agnostic features we measured in our preliminary experiments. These features include:

### v2

@@ 3.4 Causal Features Preservation
To eliminate the effect of causal features, it is essential to follow the principle of controlled experiments by keeping causal features constant while systematically manipulating spurious features. This approach isolates the impact of spurious features from that of causal features, enabling an accurate quantification of robustness against spurious features. In our framework, we introduce two methods to ensure the stability of causal features in the grounding data. Implementation details can be found in Appendix C.
@@
In general, this algorithm can be implemented by any NLI system. However, in our case, the concatenation of $G$ and $g(G)$ sometimes exceeds the context limitation of a Bert-based NLI model. Hence, we apply an LLM-based NLI system  [^4] Farquhar et al. (2024) confirms the effectiveness of the LLM-based NLI system through human annotation, demonstrating that its performance is on par with the DeBERTa-large model used in Kuhn et al. (2023). to implement the bidirectional entailment algorithm.
@@
While semantic equivalence protects causal features to the greatest extent, the perturbation may lead to the correct answer being paraphrased into an alias (e.g., "President Roosevelt" to "Roosevelt"). To address this issue, we employ a simple string-matching strategy to filter out documents that have undergone unexpected modifications.
3.5 Robustness Evaluation
[…]
$\text{RR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(C==0)$ (3)
where $N$ is the total number of instances in the dataset; $y_{i}$ and $\hat{y_{i}}$ represent the outputs of LLM for the original and perturbed instances. RR measures the proportion of instances where the RALM’s answer remains consistent (0) before and after introducing the perturbation. Similarly, WR and LR quantify the proportions of instances where the correctness of the RALM’s response changes after the perturbation, either from incorrect to correct (C==1C==-1) or from correct to incorrect ($C==1$).
Figure 3: A comparison of dataset-level metric (Acc) and instance-level metric ($R$) for robustness evaluation. ✔and ✗indicate the correctness of responses. In this example, $R$ captures instance-level unrobustness, while Acc overlooks RALMs’ sensitivity to spurious features within documents.
4 Experiments
Mistral-7B-Instruct-v0.3
Taxonomy Perturbations Known-Golden Known-Noise Unknown-Golden U-N
LR $R$ WR Org Acc LR $R$ WR Org Acc LR $R$ WR Org Acc $R$
Style Simple 7.33 85.00 7.67 73.02 73.37 4.45 91.64 3.90 10.82 10.28 7.87 82.95 9.18 56.31 57.62 98.76
Complex 6.05 87.42 6.53 73.50 3.85 92.03 4.12 11.10 6.90 85.92 7.17 56.58 98.82
[1pt/1pt] Source LLM-Generated 5.91 87.62 6.47 71.81 72.36 3.57 92.27 4.16 10.79 11.38 6.41 86.52 7.06 54.46 55.11 98.75
Self-Generated 6.30 87.06 6.64 72.15 3.94 92.02 4.04 10.89 6.26 86.80 6.94 55.14 98.77
[1pt/1pt] Logic Reverse 5.44 89.34 5.22 69.91 69.69 2.99 94.10 2.92 11.77 11.70 5.97 88.54 5.49 50.26 49.79 99.04
Random 4.47 91.87 3.66 69.10 2.43 95.15 2.42 11.76 4.18 91.44 4.38 50.46 99.27
LLM-Ranked 3.52 93.15 3.33 69.72 2.07 95.84 2.09 11.79 3.57 92.89 3.54 50.24 99.30
[1pt/1pt] Format JSON 7.96 88.53 3.51 70.81 66.35 5.15 92.68 2.17 10.98 8.00 6.95 88.92 4.13 53.32 50.50 99.02
HTML 9.30 87.03 3.67 65.18 5.89 92.36 1.74 6.83 8.36 87.39 4.25 49.22 99.01
YAML 4.75 90.90 4.35 70.41 3.88 93.24 2.87 9.97 5.05 90.53 4.42 52.69 99.06
Markdown 3.98 92.49 3.53 70.36 2.91 94.36 2.72 10.79 4.11 92.59 3.31 52.52 99.15
[1pt/1pt] Metadata Timestamp (pre) 2.62 94.90 2.48 65.04 64.90 1.28 97.61 1.11 6.83 6.66 3.15 94.45 2.40 48.08 47.33 99.67
Timestamp (post) 2.74 94.87 2.40 64.70 1.16 97.63 1.21 6.88 3.45 94.41 2.14 46.77 99.68
Datasource (wiki) 3.78 92.31 3.91 65.17 1.5 96.66 1.84 7.16 3.69 92.95 3.36 47.76 99.48
Datasource (twitter) 2.68 93.59 3.73 66.08 1.3 97.22 1.48 7.00 2.04 94.90 3.06 49.10 99.59
Table 1: Robustness evaluation results of Mistral-7B-Instruct-v0.3 on the SURE_Wiki dataset. Org indicates the accuracy on original instances, while Acc refers to the accuracy after introducing perturbations. We use Bold to mark the WR values that are higher than the LR, suggesting that the perturbation is beneficial.
In this section, we assess the robustness of RAG systems to spurious features by evaluating them on their most popular application—the Question Answering (QA) task, following the standard "retrieve-read" setting of the RAG paradigm.
4.1 Experimental Setup
Datasets
Through the steps of spurious features injection and causal features preservation, we derive the final dataset available for robustness evaluation: SURE_Wiki. The queries are drawn from the NQ-open dataset Lee et al. (2019), while our data source is English Wikipedia dump.
Models
We test two representative LLMs in our main experiments: Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct. Further implementation details are included in Appendix E.
@@ 4.1 Experimental Setup
Table 2: Statistics of the SURE_Wiki dataset for Mistral-7B-Instruct-v0.3. K-G denotes the instances composed of (Known query, Golden Document), while U-N refers to the instances consisting of (Unknown query, Noise Document). The values represents the number of instance pairs for each type of perturbations within the category of spurious features.
4.2 Experimental Results
To further analyze spurious features, we divide SURE_Wiki into four subsets based on the categories of queries and documents within each instance. A query is labeled as Known if it can be correctly answered in a closed-book setting; otherwise, it is labeled as Unknown. Documents are categorized as Golden or Noise depending on whether they contain ground truths. Notably, the distribution of the dataset is model-specific, as the classification of Known and Unknown queries is determined by the intrinsic knowledge of the target LLM. Table 2 presents dataset statistics for Mistral-7B-Instruct-v0.3, while the distribution for Llama-3.1-8B-Instruct is shown in Appendix D.
For Different Queries and Grounding Data
We report the results of Mistral-7B-Instruct and Llama-3.1-8B-Instruct in Table 1 and Table 9, respectively. For golden documents, the robustness rates of K-G and U-G are very similar for both Mistral and Llama, whereas their accuracy differ significantly. This suggests that, unlike robustness to explicit noise Wu et al. (2024b), robustness against spurious features is independent of the model’s internal prior knowledge.
When tested on noise documents, the RR remains high across all spurious features, as LLMs consistently generate incorrect responses in the absence of ground truths. Therefore, we primarily focus on the RR results for the golden documents in the following experiments.
For Different Perturbations
We observe notable differences in robustness rates across the five types of spurious features. However, within each category, the RR values for different perturbations are relatively similar. Hence, the robustness of spurious features can be estimated by averaging the RR values of their corresponding perturbations.
When further comparing perturbations within the same category, we find that while their RR values are comparable, their WR and LR can differ significantly. If the WR exceeds the LR, more instances are corrected than misanswered after introducing perturbations. This suggests that not every spurious feature is harmful and they can even be beneficial sometimes.
4.3 Further Analysis
The raw synthetic dataset is not ideal for extensive evaluation due to its large size. Furthermore, the class imbalance result in unfair comparisons across different types of spurious features. To facilitate more efficient evaluation, we extract the most challenging data from our synthetic datasets to create a lightweight benchmark: SIG_Wiki (Spurious features In Golden document from Wiki)  [^5] Specifically, we randomly select 100 instance pairs for each perturbation where both models lack robustness..
Spurious Features in Different Model Families.
To examine whether spurious features are merely artifacts of specific model choices, we evaluate a diverse set of SOTA LLMs on the SIG benchmark. The evaluated models include GPT-4O, GPT-4O-mini, Mistral-Large-Instruct  [^6] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama-3.3-70B-Instruct, Qwen2.5-72B-Instruct, and DeepSeek-V3 (671B,MoE), covering a wide range of model series and architectures. To better compare the robustness of different models, we average the RR of each perturbation within a category to derive the overall robustness for a specific type of spurious feature. The performance of six SOTA LLMs is then visualized using a radar chart, as shown in Figure 4. Despite the impressive robustness of closed-source models, they still exhibit sensitivity to certain specific perturbations. These results demonstrate that spurious features are a widespread issue across different model families, sizes, and architectures (Dense VS. MoE).
Figure 4: Robustness comparison of six SOTA LLMs.
Spurious Features Across Retrieval Settings.
To evaluate whether spurious features exist across different retrieval configurations (query sources, retrievers, and evidence sources), we additionally construct a benchmark, SIG_Trivial, following the SURE framework pipeline. This benchmark employs a production-level retriever (Bing Search) to retrieve documents from the open web, using queries from Trivial_QA Joshi et al. (2017). Further details are provided in Appendix F. Table 3 presents the average RR for each type of spurious feature. The results demonstrate that spurious features are a prevalent issue for RAG systems across different retrieval configurations.
Style Source Logic Format Meta
Mistral-7B-Instruct 88.0 94.0 94.5 94.0 99.0
with LLM-as-Judge 90.5 91.5 92.0 93.8 96.0
[1pt/1pt] Llama-3.1-8B-Inst. 87.5 93.5 93.0 90.8 97.0
with LLM-as-Judge 85.0 92.0 91.0 90.8 93.3
Table 3: Robustness evaluation of two models on SIG_Trivial, including results from LLM-as-Judge paradigm as a complementary assessment.
Reliability of Robustness Evaluations.
To validate the reliability of our string-based evaluation method, we introduce the LLM-as-Judge paradigm as a comparative assessment on SIG_Trivial benchmark. See more details in Appendix G. As illustrated in Table 3, our evaluation results similar to those of LLM-as-Judge, suggesting its effectiveness and efficiency.
4.4 Mitigating Spurious Features
Can Scaling up Model Size Solve the Problem?
To investigate the impact of parameter scale on RAG robustness, we gradually increase the size of LLM-based readers (Qwen2.5 series, ranging from 0.5B to 72B) and evaluate their robustness across five types of spurious features. As illustrated in Figure 5, the robustness rate for all spurious features shows a relatively upward trend as the model size increases. However, when we further scale the model from 32B to 72B, the RR undergoes a significant decline (except for format and meta). These findings suggest that although scaling up model size can enhance robustness to some extent, it fails to fundamentally eliminate sensitivity to spurious features.
Figure 5: Scaling analysis on Qwen2.5 series.
Are Existing Prompting Methods Effective?
We evaluate whether methods developed to improve the robustness of RALMs against explicit noise can generalize to spurious features. Previous work, such as Chain-of-Note (CON) Yu et al. (2023), aims to enhance robustness by generating thorough rationale before producing the answer. Moreover, recent breakthroughs in the reasoning capabilities of LLMs have significantly advanced the cutting edge of RAG. By integrating with reasoning models, RAG can overcome previous limitations and adapt to more complex scenarios Gao et al. (2025). Therefore, we test both CON and DeepSeel-R1 on our SIG benchmark (results shown in Table 4). Notably, the robustness rate of CON is even lower than the baseline (Qwen2.5-72B-Instruct) without applying CON. A similar phenomenon was observed in experiments with the reasoning model DeepSeek-R1 Guo et al. (2025), whose robustness was even worse than its base model, DeepSeek-V3. This indicates that the robustness against spurious features cannot be effectively improved through COT-style techniques.
Style Source Logic Format Meta
Qwen2.5-72B 78.5 76.0 88.6 92.5 95.0
+ Chain-of-Note 74.0 81.7 66.7 84.8 91.0
[1pt/1pt] DeepSeek-V3 96.5 93.6 95.6 94.0 96.5
DeepSeek-R1 84.5 87.3 83.3 87.0 87.5
Table 4: Robustness evaluation of CoN and DeepSeek-R1. Values that show improvements over the baseline are marked in bold.
Improving Robustness Using Synthetic Data.
Using the data generated by our SURE framework, we introduce two training-based mitigation strategies, Supervised Finetuning (SFT) and Direct Preference Optimization (DPO), designed to enhance the robustness of RALMs against spurious features. Specifically, we select unrobust instances, each comprising a query, a correct answer, an incorrect answer, and two golden passages (original and perturbed documents). For SFT, we train the model to consistently produce the correct answer for each query, creating two training samples by pairing the answer separately with the original and perturbed golden passages. For DPO, we train the model to prefer the correct answer over the incorrect one, likewise constructing two samples per query using both golden passages. The goal of these strategies is to teach the model to reliably generate the correct answer, regardless of whether the input documents contain spurious features. We train our methods for two epochs on a dataset of over 30k samples, and evaluate their robustness rates on the SIG_Wiki and SIG_Trivial (out-of-domain) benchmarks. The backbone model is Llama-3.1-8B-Instruct. The details of the hyperparameter settings are presented in Appendix H.
As shown in Table 5, both methods significantly enhance robustness. However, SFT performs better on in-domain data (i.e., SIG_Wiki), while DPO generalizes more effectively to out-of-domain data. These results demonstrate the effectiveness of our SURE framework for generating training data and provide strong baselines for future research.
Style Source Logic Format Meta
Llama3.1-8B (Wiki) 10.0 15.5 20.0 24.0 94.0
+ SFT 96.5 94.5 99.0 99.5 99.7
+ DPO 96.5 96.0 96.0 98.0 98.0
[1pt/1pt] Llama3.1-8B (Trivial) 87.5 93.5 93.0 90.8 97.0
+ SFT 88.5 91.5 95.0 96.3 99.0
+ DPO 94.5 94.5 97.3 95.8 98.0
Table 5: Robustness comparison of Llama-3.1-8B-Instruct trained with SFT and DPO. The upper section shows the results on SIG_Wiki, while the lower section presents the results on SIG_Trivial. The best score for each type is highlighted in bold.
5 Conclusion
In this work, we formally highlight the spurious features problem in RAG system. To quantify the robustness of RALMs against spurious features, we propose a novel evaluation framework, SURE, which includes a comprehensive taxonomy, a data synthesis pipeline, and evaluation metrics. Furthermore, leveraging the synthetic data generated by SURE, we introduce two training-based approaches that effectively improve the robustness of RALMs. Overall, our framework enables the systematic evaluation and mitigation of spurious features, paving the way for future research.
Limitations
We strive to comprehensively cover all types of spurious features that may arise in RAG scenarios. However, some unidentifiable spurious features may fall outside the scope of our taxonomy and thus fail to be quantified using the proposed SURE framework.
Ethical Considerations
All datasets used in our work are publicly released under open licenses, and all models employed for generation are open-source and licensed for research use. We strictly follow the licensing terms and usage policies of these resources.
Appendix
Appendix A Preliminary
In this section, we first define causal and spurious features in the context of RAG and then demonstrate the existence of spurious features statistically.
A.1 Causal and Spurious Features in RAG
In general, causal features are input features that have a direct causal effect on the output of predictive model Yu et al. (2020). Their relationship is rooted in causality, rather than mere statistical correlation. When it comes to Large Language Models, the meaning and intent of prompts serve as causal features that directly influence the models’ responses. In the context of RAG, causal features refer to the semantic information of grounding data.
In contrast, spurious features are input features that co-occur with causal features and are erroneously captured by the model Neuhaus et al. (2023). These features exhibit a statistical correlation with the model’s output but lack a causal relationship. Recent research has shown that LLMs are sensitive to seemingly trivial features like prompt formatting, thereby extending the definition of spurious features to LLMs Sclar et al. (2024). Similarly, we define the semantic-agnostic features of the grounding data as spurious features in RAG systems. However, conclusions drawn from in-context learning scenarios (e.g., classification and multiple-choice tasks) may not applicable to RAG scenarios, which typically involve open-ended generation tasks. Therefore, we design a preliminary experiment to validate the presence of spurious features in RAG.
A.2 Preliminary Experiment
We aim to demonstrate the semantic-agnostic features within real documents are spurious features, i.e., to reveal their impact on the output of RAG systems.
There are some challenges in revealing the influence of semantic-agnostic features. First, when retrieving from a single corpus, it is difficult to mine semantically equivalent counterparts that differ only in semantic-agnostic features. To mine appropriate documents, we introduce Contriever-msmarco, a traditional dense retriever, to recall 100 semantically similar candidates. To further eliminate the effect of causal features, documents without golden answers are filtered out, ensuring that the remaining documents have roughly consistent causal features.
Still, the differences in spurious features among these candidate documents are often minor, and their impact on model responses cannot be effectively captured using binary evaluation methods that simply judge whether an answer is correct or incorrect. Thus, more fine-grained metrics are required to detect such nuanced performance changes. Inspired by the use of LLMs as supervision signals for document utility Izacard et al. (2023); Gan et al. (2024), we introduce the oracle score, which measures fine-grained performance through calculating the log probability of generating correct answers given a specific document. The oracle score is defined as follows:
@@ A.2 Preliminary Experiment
where $x$ is the input prompt for RALMs, including the instruction $I$, grounding data $G$, and query $Q$; $y$ represents the ground truth answer; $\theta$ denotes the model parameters; and $T$ is the total length of the answer sequence  [^7] For cases with multiple answers, we compute the final score by averaging the corresponding oracle scores across all answers..
For each query, we construct document pairs by selecting the first-ranked and last-ranked candidate documents based on their oracle scores. However, the presence of various semantic-agnostic features within each document pairs makes it challenging to isolate the impact of any individual features. To assess the influence of a given feature, we compare its distribution between document sets with first- and last-ranked oracle scores. A control group is constructed by randomly sampling two document sets. If the distributions differ significantly, it suggests that RALMs are sensitive to the feature. See Appendix A for implementation details of preliminary experiments.
We test the following features: 1) Flesh Score, 2) Distinct-1, 3) Dependency Tree Depth, 4) PPL, and 5) Token Length. The results show that RALMs are sensitive to semantic-agnostic features. Nevertheless, it does not offer empirical evidence or quantitative analysis. Inspired by previous data synthesis studies Tan et al. (2024b), we use a data synthesis approach to better control feature variables and quantify the robustness of RALMs.
A.3 Preliminary Experiment Results
Using Contriever-msmarco, we recall 100 documents from the Wikipedia dump for each query in the NQ-open dataset. After filtering out documents that do not contain golden answers, we select the first-ranked and last-ranked documents based on their oracle scores for each query from the remaining documents, resulting in two sets of 2658 samples each. By comparing the differences in feature distributions between these two sets, we can assess whether RALMs exhibits sensitivity toward semantic-agnostic features. If these two sets do not belong to the same feature distribution, this can be attributed to the inherent bias of LLMs towards semantic-agnostic features. To confirm that this bias is not introduced by the dense retriever in first-stage retrieval, we establish a control group by randomly sampling two documents instead of selecting the first- and last-ranked documents.
To evaluate whether the two distributions are same, we employ the Kolmogorov-Smirnov (K-S) test. The following semantic-agnostic features are measured in the experiments: